# Distributed games with jumps: an $\alpha$-potential game Approach

### v1

@@ 2.2. NEs and $\alpha$-potential function
$|J_{i}\left(\left(u_{i}^{\prime},u_{-i}\right)\right)-J_{i}\left(\left(u_{i},u_{-i}\right)\right)-\left(\Phi\left(\left(u_{i}^{\prime},u_{-i}\right)\right)-\Phi\left(\left(u_{i},u_{-i}\right)\right)\right)|\leq\alpha.$ (2.5)
@@
Using the notion of linear derivatives, the following theorem constructs an $\alpha$-potential function for the game $\mathcal{G}$ and quantify the associate $\alpha$.
@@
$\Phi({\textbf{{u}}})=\int_{0}^{1}\sum_{j=1}^{N}\frac{\delta J_{j}}{\delta u_{j}}\left(r{\textbf{{u}}};u_{j}\right)\mathrm{d}r.$ (2.6)
@@
$\alpha\leq\frac{1}{2}\sup_{i\in[N],u^{\prime}_{i}\in\mathcal{A}_{i},u^{\prime}_{i}\in\mathcal{A}_{j},{\textbf{{u}}}\in\mathcal{A}}\sum_{j=1}^{N}\left|\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}\left({\textbf{{u}}};u_{i}^{\prime},u^{\prime\prime}_{j}\right)-\frac{\delta^{2}J_{j}}{\delta u_{j}\delta u_{i}}\left({\textbf{{u}}};u^{\prime\prime}_{j},u_{i}^{\prime}\right)\right|.$ (2.7)
[…]
@@ 3. $\alpha$-Potential Function for Distributed Games
$\frac{\delta J_{i}}{\delta u_{i}}\left({\textbf{{u}}};u_{i}^{\prime}\right)$ $=\mathbb{E}\Bigg{[}\int_{0}^{T}\begin{pmatrix}Y^{u^{\prime}_{i}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial_{x_{i}}f_{i}\\ \partial_{a_{i}}f_{i}\end{pmatrix}(t,{\textbf{X}}_{t}^{{\textbf{{u}}}},{\textbf{{u}}}_{t})\,\mathrm{d}t+(Y^{u^{\prime}_{i}}_{i,T})^{\top}(\partial_{x_{i}}g_{i})({\textbf{X}}_{T}^{{\textbf{{u}}}})\Bigg{]},$ (3.1)
@@
${\mathrm{d}}Y_{i,t}=b_{i}(t)u_{i,t}^{\prime}{\mathrm{d}}t,\quad t\in(0,T];\quad Y_{i,0}=0.$ (3.2)
@@
$\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};u^{\prime}_{i},u^{\prime\prime}_{j})$ $={\mathbb{E}}\left[\int_{0}^{T}\begin{pmatrix}Y^{u_{i}^{\prime}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial^{2}_{x_{i}x_{j}}f_{i}&\partial^{2}_{x_{i}a_{j}}f_{i}\\ \partial^{2}_{a_{i}x_{j}}f_{i}&\partial^{2}_{a_{i}a_{j}}f_{i}\end{pmatrix}(t,{\textbf{X}}_{t}^{{\textbf{{u}}}},{\textbf{{u}}}_{t})\begin{pmatrix}Y^{u_{j}^{\prime\prime}}_{j,t}\\ u^{\prime\prime}_{j,t}\end{pmatrix}{\mathrm{d}}t\right]$ (3.3)
@@
Lemma 3.1 follows directly from [6, Lemma 4.8], the convexity of $\mathcal{A}_{i}$, and the linearity of the state dynamics (2.2). The expression (3.3) of the second-order derivative $\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}$ is simpler than the formula given in [17, Equation 4.6] for general stochastic differential games, due to the fact that player $j$’s control does not affect player $i$’s state evolution.
@@
$\Phi({\textbf{{u}}})$ $=\int_{0}^{1}\sum_{i=1}^{N}\frac{\delta J_{i}}{\delta u_{i}}\left(r{\textbf{{u}}};u_{i}\right){\mathrm{d}}r$ (3.4)
@@
which depends on the aggregated behaviour of ${\textbf{Y}}^{{\textbf{{u}}}}=(Y^{u_{i}}_{i})_{i\in[N]}$ and the family of state processes $({\textbf{X}}^{r{\textbf{{u}}}})_{r\in[0,1]}$ parameterized by $r$. To further simplify the expression (3.4), the following lemma exploits the structure of the state dynamics (2.2) and (3.2), and decomposes ${\textbf{X}}^{r{\textbf{{u}}}}$ into ${\textbf{X}}^{{\textbf{{u}}}}$ and ${\textbf{Y}}^{{\textbf{{u}}}}$.
@@
The proof simply follows by noting that the process $\tilde{{\textbf{X}}}\coloneqq{\textbf{X}}^{{\textbf{{u}}}}-(1-r){\textbf{Y}}^{{\textbf{{u}}}}$ has the same initial condition and satisfies the same dynamics as ${{\textbf{X}}}^{r{\textbf{{u}}}}$.
@@
Suppose (H.1) holds. The function $\Phi:\mathcal{A}^{(N)}\to{\mathbb{R}}$ in (2.6) can be expressed as
$\Phi(\boldsymbol{u})=\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\right],$ (3.5)
@@
$\begin{split}F(t,x,y,a)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}\binom{y_{i}}{a_{i}}^{\top}\binom{\partial_{x_{i}}f_{i}}{\partial_{a_{i}}f_{i}}\left(t,x-(1-r)y,ra\right){\mathrm{d}}r,\\ G(x,y)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}y_{i}^{\top}\left(\partial_{x_{i}}g_{i}\right)\left(x-(1-r)y\right){\mathrm{d}}r.\end{split}$ (3.6)
@@
$\alpha$ $\leq\frac{1}{2}\sup_{i\in[N]}\sum_{j\in[N]\setminus\{i\}}U_{i}U_{j}\Bigg{(}TB_{i}B_{j}\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}+T^{\frac{1}{2}}B_{i}\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}+T^{\frac{1}{2}}B_{j}\|\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}$ (3.7)
@@
The expression (3.5) follows by substituting the expression ${\textbf{X}}^{r{\textbf{{u}}}}={\textbf{X}}^{{\textbf{{u}}}}-(1-r){\textbf{Y}}^{{\textbf{{u}}}}$ into Lemma 3.2, and applying Fubini’s theorem.
@@
$\begin{split}&\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};u^{\prime}_{i},u^{\prime\prime}_{j})-\frac{\delta^{2}J_{j}}{\delta u_{j}\delta u_{i}}({\textbf{{u}}};u^{\prime\prime}_{j},u^{\prime}_{i})\\ &={\mathbb{E}}\left[\int_{0}^{T}\begin{pmatrix}Y^{u_{i}^{\prime}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}&\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\\ \partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}&\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j}\end{pmatrix}(t,\cdot)\begin{pmatrix}Y^{u_{j}^{\prime\prime}}_{j,t}\\ u^{\prime\prime}_{j,t}\end{pmatrix}{\mathrm{d}}t+(Y^{u_{i}^{\prime}}_{i,T})^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j})({\textbf{X}}^{{\textbf{{u}}}}_{T})Y^{u_{j}^{\prime\prime}}_{j,T}\right],\end{split}$ (3.8)
@@
$\mathbb{E}\left[\left|Y_{i,t}^{u_{i}^{\prime}}\right|^{2}\right]$ $=\mathbb{E}\left[\left|\int_{0}^{t}b_{i}(v)u_{i}^{\prime}(v){\mathrm{d}}v\right|^{2}\right]\leq\mathbb{E}\left[\int_{0}^{t}|b_{i}(v)|^{2}{\mathrm{d}}v\int_{0}^{t}|u_{i}^{\prime}(v)|^{2}{\mathrm{d}}v\right]=\|b_{i}\|_{L^{2}}^{2}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}^{2}.$ (3.9)
@@
$\begin{split}&\left|{\mathbb{E}}\left[\int_{0}^{T}\left(Y^{u^{\prime}_{i}}_{i,t}\right)^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)Y^{u^{\prime\prime}_{j}}_{j,t}{\mathrm{d}}t\right]\right|\leq{\mathbb{E}}\left[\int_{0}^{T}|Y^{u^{\prime}_{i}}_{i,t}|\|(\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)\|_{\textrm{sp}}|Y^{u^{\prime\prime}_{j}}_{j,t}|{\mathrm{d}}t\right]\\ &\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}{\mathbb{E}}\left[\int_{0}^{T}|Y^{u^{\prime}_{i}}_{i,t}||Y^{u^{\prime\prime}_{j}}_{j,t}|{\mathrm{d}}t\right]\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|Y^{u^{\prime}_{i}}_{i}\|_{\mathcal{H}^{2}}\|Y^{u^{\prime\prime}_{j}}_{j}\|_{\mathcal{H}^{2}}\\ &\leq T\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{i}\|_{L^{2}}\|b_{j}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}.\end{split}$ (3.10)
@@
$\begin{split}&\Bigg{|}{\mathbb{E}}\Bigg{[}\int_{0}^{T}\left(Y^{u^{\prime}_{i}}_{i,t}\right)^{\top}(\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j})(t,\cdot)u_{j,t}^{\prime\prime}{\mathrm{d}}t\Bigg{]}\Bigg{|}\leq\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|Y^{u^{\prime}_{i}}_{i}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}\\ &\leq T^{\frac{1}{2}}\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{i}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}},\end{split}$ (3.11)
@@
$\begin{split}\left|{\mathbb{E}}\left[\int_{0}^{T}\left(u^{\prime}_{i,t}\right)^{\top}(\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)Y_{j,t}^{u^{\prime\prime}_{j}}{\mathrm{d}}t\right]\right|&\leq T^{\frac{1}{2}}\|\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{j}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}},\\ \left|{\mathbb{E}}\left[\int_{0}^{T}\left(u^{\prime}_{i,t}\right)^{\top}(\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j})(t,\cdot)u_{j,t}^{\prime\prime}{\mathrm{d}}t\right]\right|&\leq\|\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}.\end{split}$ (3.12)
@@
${\mathbb{E}}\left[(Y^{u_{i}^{\prime}}_{i,T})^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j})({\textbf{X}}^{{\textbf{{u}}}}_{T})Y^{u_{j}^{\prime\prime}}_{j,T}\right]$ $\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j}\|_{L^{\infty}}{\mathbb{E}}\left[|Y^{u_{i}^{\prime}}_{i,T}||Y^{u_{j}^{\prime\prime}}_{j,T}|\right]$ (3.13)
@@
$\partial^{2}_{x_{i}x_{j}}f_{i}=\partial^{2}_{x_{i}x_{j}}f_{j},\quad\partial^{2}_{a_{i}x_{j}}f_{i}=\partial^{2}_{a_{i}x_{j}}f_{j},\quad\partial^{2}_{a_{i}a_{j}}f_{i}=\partial^{2}_{a_{i}a_{j}}f_{j},\quad\partial^{2}_{x_{i}x_{j}}g_{i}=\partial^{2}_{x_{i}x_{j}}g_{j}.$ (3.14)
@@
$\bar{\Phi}({\textbf{{u}}})\coloneqq\mathbb{E}\left[\int_{0}^{T}\bar{F}(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+\bar{G}({\textbf{X}}^{{\textbf{{u}}}}_{T})\right],$ (3.15)
@@
$\bar{\Phi}({\textbf{{u}}})=\Phi({\textbf{{u}}})+\mathbb{E}\left[\int_{0}^{T}\bar{F}(t,{\textbf{X}}^{\boldsymbol{0}}_{t},\boldsymbol{0}){\mathrm{d}}t+\bar{G}({\textbf{X}}^{\boldsymbol{0}}_{T})\right],$ (3.16)
@@
Under the symmetry condition (3.14), the fact that the function $\bar{\Phi}$ in (3.15) is a potential function for the game $\mathcal{G}$ follows from analogous arguments to those used for distributed games with Markov policies in [18, Theorem 3.2]. Since both $\bar{\Phi}$ and $\Phi$ are potential functions for the game $\mathcal{G}$, it holds for all $u\in\mathcal{A}$,

### v2

@@ 2.2. NEs and $\alpha$-potential function
(2.5) $|J_{i}\left(\left(u_{i}^{\prime},u_{-i}\right)\right)-J_{i}\left(\left(u_{i},u_{-i}\right)\right)-\left(\Phi\left(\left(u_{i}^{\prime},u_{-i}\right)\right)-\Phi\left(\left(u_{i},u_{-i}\right)\right)\right)|\leq\alpha.$
@@
Using the notion of linear derivatives, the following theorem constructs an $\alpha$-potential function for the game $\mathcal{G}$ and quantifies the associated $\alpha$.
@@
(2.6) $\Phi({\textbf{{u}}})=\int_{0}^{1}\sum_{j=1}^{N}\frac{\delta J_{j}}{\delta u_{j}}\left(r{\textbf{{u}}};u_{j}\right)\mathrm{d}r.$
@@
(2.7) $\alpha\leq\sup_{i\in[N],u^{\prime}_{i}\in\mathcal{A}_{i},{\color[rgb]{0,0,0}u_{j}^{\prime\prime}}\in\mathcal{A}_{j},{\textbf{{u}}}\in\mathcal{A}}\sum_{j=1}^{N}\left|\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}\left({\textbf{{u}}};u_{i}^{\prime},u^{\prime\prime}_{j}\right)-\frac{\delta^{2}J_{j}}{\delta u_{j}\delta u_{i}}\left({\textbf{{u}}};u^{\prime\prime}_{j},u_{i}^{\prime}\right)\right|.$
Proof.
Proposition 2.1 follows as a special case of [17, Theorem 2.5], using the specific choice ${\textbf{{z}}}=0$. With this choice, the bound in (2.7) is tighter than the general upper bound on $\alpha$ provided in [17, Equation 1.3], as it no longer involves a multiplicative constant 22 used in [17]. ∎
@@ 3. $\alpha$-Potential Function for Distributed Games
(3.1) $\frac{\delta J_{i}}{\delta u_{i}}\left({\textbf{{u}}};u_{i}^{\prime}\right)$ $=\mathbb{E}\Bigg[\int_{0}^{T}\begin{pmatrix}Y^{u^{\prime}_{i}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial_{x_{i}}f_{i}\\ \partial_{a_{i}}f_{i}\end{pmatrix}(t,{\textbf{X}}_{t}^{{\textbf{{u}}}},{\textbf{{u}}}_{t})\,\mathrm{d}t+(Y^{u^{\prime}_{i}}_{i,T})^{\top}(\partial_{x_{i}}g_{i})({\textbf{X}}_{T}^{{\textbf{{u}}}})\Bigg],$
@@
(3.2) ${\mathrm{d}}Y_{i,t}=b_{i}(t)u_{i,t}^{\prime}{\mathrm{d}}t,\quad t\in(0,T];\quad Y_{i,0}=0.$
@@
(3.3) $\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};u^{\prime}_{i},u^{\prime\prime}_{j})$ $={\mathbb{E}}\left[\int_{0}^{T}\begin{pmatrix}Y^{u_{i}^{\prime}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial^{2}_{x_{i}x_{j}}f_{i}&\partial^{2}_{x_{i}a_{j}}f_{i}\\ \partial^{2}_{a_{i}x_{j}}f_{i}&\partial^{2}_{a_{i}a_{j}}f_{i}\end{pmatrix}(t,{\textbf{X}}_{t}^{{\textbf{{u}}}},{\textbf{{u}}}_{t})\begin{pmatrix}Y^{u_{j}^{\prime\prime}}_{j,t}\\ u^{\prime\prime}_{j,t}\end{pmatrix}{\mathrm{d}}t\right]$
@@
Proof.
The proof follows directly from [8, Lemma 4.8], the convexity of $\mathcal{A}_{i}$, and the linearity of the state dynamics (2.2). ∎
The expression (3.3) of the second-order derivative $\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}$ is simpler than the formula given in [17, Equation 4.6] for general stochastic differential games, due to the fact that player $j$’s control does not affect player $i$’s state evolution.
@@
(3.4) $\Phi({\textbf{{u}}})$ $=\int_{0}^{1}\sum_{i=1}^{N}\frac{\delta J_{i}}{\delta u_{i}}\left(r{\textbf{{u}}};u_{i}\right){\mathrm{d}}r$
@@
which depends on the aggregated behavior of the processes $({\textbf{X}}^{r{\textbf{{u}}}})_{r\in[0,1]}$ parameterized by $r$.
To simplify the expression (3.4), the following lemma establishes a separation principle, which exploits the linearity of the dynamics (2.2) and (3.2), and decomposes ${\textbf{X}}^{r{\textbf{{u}}}}$ into ${\textbf{X}}^{{\textbf{{u}}}}$ and ${\textbf{Y}}^{{\textbf{{u}}}}$.
@@
Proof.
The proof simply follows by noting that the process $\tilde{{\textbf{X}}}\coloneqq{\textbf{X}}^{{\textbf{{u}}}}-(1-r){\textbf{Y}}^{{\textbf{{u}}}}$ has the same initial condition and satisfies the same dynamics as ${{\textbf{X}}}^{r{\textbf{{u}}}}$. ∎
@@
Suppose (H.1) holds. The function $\Phi:\mathcal{A}\to{\mathbb{R}}$ in (2.6) can be expressed as
(3.5) $\Phi(\boldsymbol{u})=\mathbb{E}$ $\left[\int_{0}^{T}F(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{Y}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+G({\textbf{X}}^{{\textbf{{u}}}}_{T},{\textbf{Y}}^{{\textbf{{u}}}}_{T})\right],$
@@
(3.6) $\begin{split}F(t,x,y,a)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}\binom{y_{i}}{a_{i}}^{\top}\binom{\partial_{x_{i}}f_{i}}{\partial_{a_{i}}f_{i}}\left(t,x-(1-r)y,ra\right){\mathrm{d}}r,\\ G(x,y)&\coloneqq\sum_{i=1}^{N}\int_{0}^{1}y_{i}^{\top}\left(\partial_{x_{i}}g_{i}\right)\left(x-(1-r)y\right){\mathrm{d}}r.\end{split}$
@@
(3.7) $\alpha$ $\leq\sup_{i\in[N]}\sum_{j\in[N]\setminus\{i\}}U_{i}U_{j}\Bigg(TB_{i}B_{j}\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}+T^{\frac{1}{2}}B_{i}\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}+T^{\frac{1}{2}}B_{j}\|\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}$
@@
The expression (3.5) follows by substituting the expression ${\textbf{X}}^{r{\textbf{{u}}}}={\textbf{X}}^{{\textbf{{u}}}}-(1-r){\textbf{Y}}^{{\textbf{{u}}}}$ from Lemma 3.2 into (3.4), and applying Fubini’s theorem.
@@
(3.8) $\begin{split}&\frac{\delta^{2}J_{i}}{\delta u_{i}\delta u_{j}}({\textbf{{u}}};u^{\prime}_{i},u^{\prime\prime}_{j})-\frac{\delta^{2}J_{j}}{\delta u_{j}\delta u_{i}}({\textbf{{u}}};u^{\prime\prime}_{j},u^{\prime}_{i})\\ &={\mathbb{E}}\left[\int_{0}^{T}\begin{pmatrix}Y^{u_{i}^{\prime}}_{i,t}\\ u_{i,t}^{\prime}\end{pmatrix}^{\top}\begin{pmatrix}\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}&\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\\ \partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}&\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j}\end{pmatrix}(t,\cdot)\begin{pmatrix}Y^{u_{j}^{\prime\prime}}_{j,t}\\ u^{\prime\prime}_{j,t}\end{pmatrix}{\mathrm{d}}t+(Y^{u_{i}^{\prime}}_{i,T})^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j})({\textbf{X}}^{{\textbf{{u}}}}_{T})Y^{u_{j}^{\prime\prime}}_{j,T}\right],\end{split}$
@@
(3.9) $\mathbb{E}\left[\left|Y_{i,t}^{u_{i}^{\prime}}\right|^{2}\right]$ $=\mathbb{E}\left[\left|\int_{0}^{t}b_{i}(v)u_{i}^{\prime}(v){\mathrm{d}}v\right|^{2}\right]\leq\mathbb{E}\left[\int_{0}^{t}|b_{i}(v)|^{2}{\mathrm{d}}v\int_{0}^{t}|u_{i}^{\prime}(v)|^{2}{\mathrm{d}}v\right]=\|b_{i}\|_{L^{2}}^{2}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}^{2}.$
@@
(3.10) $\begin{split}&\left|{\mathbb{E}}\left[\int_{0}^{T}\left(Y^{u^{\prime}_{i}}_{i,t}\right)^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)Y^{u^{\prime\prime}_{j}}_{j,t}{\mathrm{d}}t\right]\right|\leq{\mathbb{E}}\left[\int_{0}^{T}|Y^{u^{\prime}_{i}}_{i,t}|\|(\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)\|_{\textrm{sp}}|Y^{u^{\prime\prime}_{j}}_{j,t}|{\mathrm{d}}t\right]\\ &\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}{\mathbb{E}}\left[\int_{0}^{T}|Y^{u^{\prime}_{i}}_{i,t}||Y^{u^{\prime\prime}_{j}}_{j,t}|{\mathrm{d}}t\right]\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|Y^{u^{\prime}_{i}}_{i}\|_{\mathcal{H}^{2}}\|Y^{u^{\prime\prime}_{j}}_{j}\|_{\mathcal{H}^{2}}\\ &\leq T\|\partial^{2}_{x_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{i}\|_{L^{2}}\|b_{j}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}.\end{split}$
@@
(3.11) $\begin{split}&\Bigg|{\mathbb{E}}\Bigg[\int_{0}^{T}\left(Y^{u^{\prime}_{i}}_{i,t}\right)^{\top}(\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j})(t,\cdot)u_{j,t}^{\prime\prime}{\mathrm{d}}t\Bigg]\Bigg|\leq\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|Y^{u^{\prime}_{i}}_{i}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}\\ &\leq T^{\frac{1}{2}}\|\partial^{2}_{x_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{i}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}},\end{split}$
@@
(3.12) $\begin{split}\left|{\mathbb{E}}\left[\int_{0}^{T}\left(u^{\prime}_{i,t}\right)^{\top}(\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j})(t,\cdot)Y_{j,t}^{u^{\prime\prime}_{j}}{\mathrm{d}}t\right]\right|&\leq T^{\frac{1}{2}}\|\partial^{2}_{a_{i}x_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|b_{j}\|_{L^{2}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}},\\ \left|{\mathbb{E}}\left[\int_{0}^{T}\left(u^{\prime}_{i,t}\right)^{\top}(\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j})(t,\cdot)u_{j,t}^{\prime\prime}{\mathrm{d}}t\right]\right|&\leq\|\partial^{2}_{a_{i}a_{j}}\Delta^{f}_{i,j}\|_{L^{\infty}}\|u_{i}^{\prime}\|_{\mathcal{H}^{2}}\|u_{j}^{\prime\prime}\|_{\mathcal{H}^{2}}.\end{split}$
@@
(3.13) ${\mathbb{E}}\left[(Y^{u_{i}^{\prime}}_{i,T})^{\top}(\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j})({\textbf{X}}^{{\textbf{{u}}}}_{T})Y^{u_{j}^{\prime\prime}}_{j,T}\right]$ $\leq\|\partial^{2}_{x_{i}x_{j}}\Delta^{g}_{i,j}\|_{L^{\infty}}{\mathbb{E}}\left[|Y^{u_{i}^{\prime}}_{i,T}||Y^{u_{j}^{\prime\prime}}_{j,T}|\right]$
@@
(3.14) $\partial^{2}_{x_{i}x_{j}}f_{i}=\partial^{2}_{x_{i}x_{j}}f_{j},\quad\partial^{2}_{a_{i}x_{j}}f_{i}=\partial^{2}_{a_{i}x_{j}}f_{j},\quad\partial^{2}_{a_{i}a_{j}}f_{i}=\partial^{2}_{a_{i}a_{j}}f_{j},\quad\partial^{2}_{x_{i}x_{j}}g_{i}=\partial^{2}_{x_{i}x_{j}}g_{j}.$
@@
(3.15) $\bar{\Phi}({\textbf{{u}}})\coloneqq\mathbb{E}\left[\int_{0}^{T}\bar{F}(t,{\textbf{X}}^{{\textbf{{u}}}}_{t},{\textbf{{u}}}_{t}){\mathrm{d}}t+\bar{G}({\textbf{X}}^{{\textbf{{u}}}}_{T})\right],$
@@
(3.16) $\bar{\Phi}({\textbf{{u}}})=\Phi({\textbf{{u}}})+\mathbb{E}\left[\int_{0}^{T}\bar{F}(t,{\textbf{X}}^{\boldsymbol{0}}_{t},\boldsymbol{0}){\mathrm{d}}t+\bar{G}({\textbf{X}}^{\boldsymbol{0}}_{T})\right],$
@@
Under the symmetry condition (3.14), the fact that the function $\bar{\Phi}$ in (3.15) is a potential function for the game $\mathcal{G}$ follows from analogous arguments to those used for distributed games with Markov policies in [20, Theorem 3.2]. Since both $\bar{\Phi}$ and $\Phi$ are potential functions for the game $\mathcal{G}$, it holds for all $u\in\mathcal{A}$,